# Datasets for Fairness in Language Models: An In-Depth Survey

### v1

@@ 3.3.4 HolisticBias
Dataset Taxonomy. HolisticBias belongs to the class of prompt-based bias datasets. It generates naturalistic or semi-naturalistic sentences by inserting identity descriptors and noun phrases into 26 hand-crafted sentence templates. These templates include statements, questions, and descriptions that vary across grammatical structures (e.g., declaratives, interrogatives, first- and third-person frames). Example prompts might include conversational frames such as “I’m sorry to hear that”, immediately following an identity descriptor to elicit subtle bias. The dataset was constructed using community-informed descriptors to ensure inclusiveness and ethical representation. The result is a large corpus of approximately 472,000 English prompts, generated via combinatorial expansion of descriptors, 30 common noun phrases, and sentence templates. While most descriptors correspond to single demographic axes, some intersecting terms like “Black woman” also appear. HolisticBias is publicly available through HuggingFace Datasets under a CC BY-SA license, with a multilingual extension (MMHB) that translates descriptors into more than 50 languages.
@@
Strengths and Limitations. HolisticBias’s major strength lies in its comprehensiveness and flexibility. It supports a variety of evaluation protocols, including token probability analysis, generation bias quantification, and classification robustness. Its participatory design process enhances the quality and cultural sensitivity of the descriptors. The multilingual extension (MMHB) further enables fairness studies across languages and cultural contexts. However, some prompts—despite careful template design—may sound unnatural in real-world conversations, potentially affecting downstream evaluations. Additionally, because the dataset primarily centers on explicit identity mentions in isolated sentences, it may fall short in capturing context-dependent or implicit bias. Its sheer size, while enabling breadth, can also introduce computational burdens for full-scale evaluation. Finally, evolving identity terminology and overlapping taxonomies (e.g., “Jewish” as both religious and ethnic) present ongoing classification challenges.
@@
Representativeness Bias. We evaluate representativeness bias in the HolisticBias dataset using the KL divergence-based formulation defined in Section 3.3, which measures the divergence between the empirical distribution $P_{\mathcal{D}}(A)$ and a reference population distribution $P_{\mathcal{P}}(A)$. This method allows us to quantify how well the dataset reflects real-world demographic proportions along protected axes such as race, religion, and gender. Reference distributions were drawn from U.S. Bureau of Labor Statistics (2024) and Pew Research datasets. To bridge the gap between the dataset’s fine-grained identity descriptors (e.g., “Buddhist”, “Afro-Latine”) and high-level reference categories (e.g., non-Christian, Black), we implemented semantic mapping dictionaries across multiple axes, including race/ethnicity, religion, gender/sex, and sexual orientation. This approach enabled aggregation of 14,346 religion-related samples, 1,594 gender samples, and 8,550 race/ethnicity samples into population-level categories for KL divergence computation.
@@
Annotation Bias. In the context of HolisticBias, direct analysis of annotation bias is not feasible because the dataset does not provide gold labels such as toxicity, sentiment, or other human- or model-generated annotations. Instead, HolisticBias is designed as a prompt-based resource for probing model fairness, with each instance consisting of a sentence template and a demographic descriptor, but lacking any associated ground-truth label. As a result, the dataset itself is neutral with respect to annotation bias: any observed annotation bias would necessarily arise from the external labeling function (e.g., a toxicity classifier or human raters) applied to the dataset, rather than from the dataset’s construction. To empirically assess annotation bias using HolisticBias, one must first annotate the dataset with a consistent labeling function and then compare the expected label values across demographic groups as formalized in Equation 2. This approach enables researchers to disentangle biases introduced by annotation instruments from those inherent in the dataset’s representativeness or design.
Therefore, HolisticBias is not suitable for stereotype leakage analysis via PMI or MI, as it lacks the spontaneous co-occurrence structure necessary for these metrics. This contrasts with naturalistic corpora such as Wikipedia or Reddit, where co-occurrence patterns are organically present and stereotype leakage can be meaningfully quantified. HolisticBias’s strength lies in its ability to test axis-specific model behavior under tightly controlled inputs, not in revealing embedded societal biases within its own textual structure.
@@
Differential Metric Bias. We probe HolisticBias with four surrogate scorers—Gender Polarity, VADER Sentiment, RoBERTa—Regard, and Perspective–API Toxicity. Only the toxicity-related graphs are shown here; all other metric plots are included in Appendix A.
(a) Gender Polarity. The Gender Polarity metric (see Appendix A) measures token-level association of words with a male/female polarity axis.
  • The heatmap (Fig. 44) further reveals specific descriptors aligned with male or female polarity.
(b) Sentiment. The VADER Sentiment metric evaluates sentence-level polarity. Axis-level differences are minimal (Kruskal–Wallis $p=0.48$), and descriptor-level variation (Appendix A) is often tied to emotionally charged words (e.g., “curvy”, “bankrupt”).
(c) Regard. RoBERTa-based Regard scores (Appendix A) highlight positive/negative connotations. Despite uniformly high medians, individual terms like “drug-addicted” (low) and “Christian” (high) introduce descriptor-level variance.
(d) Toxicity and Sub-Dimensions. We evaluate differential metric bias in the HolisticBias dataset using the Perspective API, which provides six sub-dimensions: TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, PROFANITY, and THREAT. These metrics enable us to capture nuanced differences in how identity-related prompts are perceived in terms of harmful or offensive content.
@@
This pattern persists across sub-dimensions as well. IDENTITY_ATTACK scores often co-occur with TOXICITY patterns, particularly for racial and religious descriptors (see Appendix A.4, Figures 52 and 53). Likewise, SEVERE_TOXICITY and THREAT scores are relatively low overall but spike for marginalized groups when prompts contain specific identity terms or stigmatized descriptors. INSULT scores also align closely with descriptors related to disability, age, or appearance, indicating a tendency for models to associate these identities with ridicule or demeaning language (see Appendix A.4, Figures 54 and 55).
@@
3.4 Sentiment and Classification Bias Datasets
This subsection discusses datasets created to evaluate fairness specifically in the context of downstream NLP tasks like sentiment analysis and text classification. These datasets typically use controlled templates or minimal pairs where demographic terms are varied within sentences relevant to the classification task, allowing assessment of whether identity attributes unfairly influence model predictions.
@@ 3.4.1 Equity Evaluation Corpus (EEC)
Description. The Equity Evaluation Corpus (EEC), introduced by Kiritchenko and Mohammad (2018), is a benchmark dataset designed to detect demographic bias—specifically race and gender bias—in sentiment analysis systems. It consists of sentence templates with controlled lexical substitutions where only a race- or gender-associated term (e.g., names, pronouns) is varied, enabling a counterfactual comparison of sentiment intensity scores. The goal is to identify whether systems yield different sentiment predictions for sentences that are semantically identical except for demographic references.
Dataset Taxonomy. EEC falls under the category of sentiment and classification bias datasets. Its structure is based on a minimal-pair counterfactual design using templates with emotion-laden contexts. Sentence variants are created by swapping gender- or race-associated words, allowing isolated evaluation of bias effects. The dataset is monolingual (English), publicly available, and associated with a *SEM 2018 paper. It focuses on sentiment-related biases in text classification systems and is likely distributed under a permissive license.
Intrinsic Characteristics. The dataset comprises 8,640 English sentences generated from a relatively small number of core templates. Each sentence includes emotional expressions, and variants differ only in the demographic term used. The text-based format supports standardized input to sentiment analysis models, making it straightforward for quantitative bias evaluations.
@@
However, the dataset has notable limitations. It is highly sensitive to minor rewordings of template structures, raising concerns about robustness. Its reliance on binary gender terms and a small number of templates limits both its inclusivity and linguistic diversity. Moreover, the template-based nature of the data leads to concerns about artificiality and limited ecological validity. Even the dataset creators acknowledge that EEC should be viewed as a complementary tool rather than a comprehensive benchmark for bias.
Bias Analysis. Using the methodology of 2.3, we obtain:
@@
  • Gender. The corpus is split exactly 50 % male / 50 % female. Using 2024 U.S. Bureau of Labor Statistics (BLS) workforce figures (51.2 % male / 48.8 % female) as the reference, the KL divergence is negligible ($D_{\mathrm{KL}}{=}0.0003$), confirming that gender representativeness is virtually perfect.
[…]
Annotation Bias. The Equity Evaluation Corpus (EEC) is constructed using a set of predefined templates and identity variables, with each sentence’s label (emotion category) determined deterministically by the template used. As a result, the EEC does not contain human- or model-generated annotation scores or gold labels that could reflect subjective judgment. Consequently, the dataset itself is not susceptible to annotation bias, as defined by systematic discrepancies in ground-truth labels across demographic groups. For any pair of sentences differing only in the protected attribute (e.g., gender or race), the assigned label remains constant by design. Therefore, the annotation bias metric for the EEC dataset is identically zero. However, the EEC is frequently employed to evaluate annotation bias in external systems—such as sentiment or toxicity classifiers—by measuring whether these systems assign different scores to minimal pairs that differ only in identity terms. In this context, the EEC serves as a valuable instrument for detecting annotation bias in models, even though the dataset itself is free from such bias.
and the corpus-level mutual information
$\text{MI}(W_{G};W_{T})\approx 0.$ (30)
Differential Metric Bias. This type of bias arises when model outputs exhibit systematically different values across demographic groups (e.g., gender, race) on standard evaluation metrics. To measure this, we analyze four commonly used diagnostic dimensions: Gender Polarity, Sentiment, Regard, and Toxicity, across emotion, gender, and race groups. Results are visualized in Figures 21, 22, 23and 24.
@@
    Gender Polarity: While unigram scores remain near-zero, indicating no explicit word-level gender bias, both the gender_max_score and gender_wavg_score show noticeable differences. Female-associated samples tend to have significantly higher polarity values than male-associated ones. For example, across emotions, sadness and joy show slightly higher polarity variance compared to anger and fear. Race-wise, European-associated samples show greater polarity spread than African-American counterparts. These results suggest subtle alignment differences between generated text and gendered semantic directions in the embedding space.
@@
[…]
@@
[…]
@@
    Toxicity: As expected for a balanced template corpus, overall toxicity scores are low. Emotion is the dominant factor (anger $>$ fear $\approx$ sadness $>$ joy), but no reliable demographic gap is detected:
@@
[…]
@@
    Figure 25: Differential Metric Bias: Toxicity score distributions (toxicity, severe toxicity, identity attack, insult, threat) by Emotion, Gender, and Race using Perspective API.
Because EEC sentences come in counterfactual minimal pairs, any statistically significant score gap must arise from the metric itself. Gender Polarity exhibits the strongest instrument bias, Sentiment a moderate bias, Regard a weak bias, and Toxicity none. Therefore, researchers must normalise or otherwise correct these scorers before attributing unfairness to models evaluated on EEC.
@@ 3.4.2 Bias NLI
Description. Bias NLI refers to a collection of datasets and methodological efforts designed to probe bias—particularly gender bias—within the Natural Language Inference (NLI) task. As a classification task where models predict entailment, contradiction, or neutrality between a premise and a hypothesis, NLI serves as a valuable testbed for investigating whether models make predictions based on logical inference or rely on spurious correlations. Many approaches introduce template-based counterfactuals by varying demographic terms in hypotheses while preserving the ground truth label. Others analyze existing NLI datasets, such as SNLI, to reveal annotation artifacts or demographic term dependencies that may reinforce social biases.
Dataset Taxonomy. Bias NLI falls under the category of classification bias datasets, with a strong emphasis on sentiment and stereotype-related evaluation. Its structure varies from manually curated hypothesis–premise pairs to large-scale template-generated corpora, some exceeding five million examples. These datasets are primarily in English, and their construction draws either from transforming existing corpora or creating synthetic sentence pairs to test model sensitivity to demographic cues. The typology of bias explored in Bias NLI primarily involves gender, though race and religion have also been considered. In addition to social bias, these datasets shed light on structural issues such as hypothesis-only bias, contradiction-word bias, and the more recently identified attestation bias, where models depend on memorized propositions rather than performing genuine inference. Accessibility varies depending on the particular dataset, but many are publicly released alongside research publications.
@@
Strengths and Limitations. A major strength of Bias NLI is its ability to isolate whether models are truly performing inference or instead exploiting surface-level cues linked to protected attributes. The diagnostic nature of these datasets allows researchers to pinpoint specific sources of bias, such as hypothesis-only bias or word-overlap effects. However, the reliance on template-generated sentences can introduce artificiality and make models overly sensitive to phrasing. In some cases, evaluations using Bias NLI may reflect known NLI dataset artifacts more than the presence of social bias, complicating interpretation. Moreover, template-based construction may exacerbate attestation bias, causing models to memorize and reuse stereotypical associations. Finally, while these datasets are effective at highlighting known issues, they may underexplore other forms of latent bias or intersectional effects.
@@
Representativeness Bias. We concatenated the train, dev, and test splits of SNLI ($N{=}570{,}152$ premise–hypothesis pairs) and extracted gender information by keyword matching over both fields.Pronouns and role nouns such as he, she, man, woman, father, mother, boy, girl, etc. were used. Gender could be detected in $434{,}367$ instances ($76.2\%$ coverage). The resulting empirical distribution is
@@
whereas the 2024 U.S. Bureau of Labor Statistics baseline is
@@
Applying the KL-divergence metric defined in Section  2.3,
@@
reveals a substantial $27$-percentage-point over-representation of masculine references. This skew implies that fairness evaluations based on BiasNLI may inflate performance on male-coded language while offering weaker evidence of model behaviour on female-coded inputs, underscoring the need for rebalancing or stratified reporting.
Annotation Bias. BiasNLI pairs every premise with three template-generated hypotheses, explicitly assigned the labels entailment, neutral, and contradiction. Because these labels are deterministic by construction—not crowd-annotated—the expected label distribution is identical for every protected attribute. Hence annotation bias is not a meaningful diagnostic for this dataset; any discrepancy would merely reflect the dataset’s fixed three-way design rather than systematic rater bias.
Stereotype Leakage. To measure how demographic references co-occur with socially stereotyped traits, we apply the PMI and MI estimators introduced in Section 3. Group and trait vocabularies—comprising gender, race, religion, age, and disability markers, as well as traits related to occupation, personality, behavior, and social status—are pre-defined and used throughout the analysis. For each sentence in the BiasNLI corpus, we slide a token window of size 5 and count co-occurrences of $(w_{g},w_{t})$ pairs. Co-occurrence probabilities are smoothed and normalized to compute both pairwise PMI and corpus-level MI.
@@
Figure 26 illustrates the Top-10 most stereotypical co-occurrence pairs according to PMI. Among the highest are (jewish, rich), (black, unsuccessful), (him, aggressive), and (african, poverty). These combinations suggest stereotypical associations between religion and wealth, race and socioeconomic failure, and gender and emotional dominance—aligning with historically attested social biases.

### v2

@@ 3.3.4 HolisticBias
Dataset Taxonomy. HolisticBias is classified as an open-ended evaluation dataset, requiring models to generate text freely in response to prompts without predefined output constraints. It employs a participatory framework that systematically spans 13 demographic axes using nearly 600 identity descriptors, generating approximately 472,000 English prompts through combinatorial expansion of descriptors, noun phrases, and sentence templates. The dataset’s flexible design enables comprehensive bias evaluation across a wide spectrum of identities while maintaining experimental control through its template-based construction.
In terms of source, HolisticBias employs template-based construction combined with community-informed descriptors. The dataset generates naturalistic or semi-naturalistic sentences by inserting identity descriptors and noun phrases into 26 hand-crafted sentence templates, ensuring inclusiveness and ethical representation through community input while maintaining experimental control over sentence variability.
The dataset exhibits multilingual linguistic coverage, with the original version focusing on English and a multilingual extension (MMHB) that translates descriptors into more than 50 languages. This multilingual approach enables cross-cultural fairness evaluations and helps identify language-specific biases, addressing the limitations of English-centric datasets.
Regarding bias typology, HolisticBias targets multiple demographic characteristic biases with comprehensive coverage including intersectional categories. While most descriptors correspond to single demographic axes, some intersecting terms like “Black woman” also appear, providing more nuanced evaluation of bias across multiple protected attributes simultaneously.
Finally, HolisticBias demonstrates high accessibility as a publicly available dataset through HuggingFace Datasets under a CC BY-SA license. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
@@
Strengths and Limitations. HolisticBias’s major strength lies in its comprehensiveness and flexibility. It supports a variety of evaluation protocols, including token probability analysis, generation bias quantification, and classification robustness. Its participatory design process enhances the quality and cultural sensitivity of the descriptors. The multilingual extension (MMHB) further enables fairness studies across languages and cultural contexts. However, some prompts, despite careful template design, may sound unnatural in real-world conversations, potentially affecting downstream evaluations. Additionally, because the dataset primarily centers on explicit identity mentions in isolated sentences, it may fall short in capturing context-dependent or implicit bias. Its sheer size, while enabling breadth, can also introduce computational burdens for full-scale evaluation. Finally, evolving identity terminology and overlapping taxonomies (e.g., “Jewish” as both religious and ethnic) present ongoing classification challenges.
@@
Representativeness Bias. We evaluate representativeness bias in the HolisticBias dataset using the KL divergence-based formulation defined in Section 3.3, which measures the divergence between the empirical distribution $P_{\mathcal{D}}(A)$ and a reference population distribution $P_{\mathcal{P}}(A)$. This method allows us to quantify how well the dataset reflects real-world demographic proportions along protected axes such as race, religion, and gender. Reference distributions were drawn from the 2020 Census (for race, gender, and age), 2022 American Community Survey (for nationality and disability), Pew Research Center 2021 (for religion and income), Gallup 2023 (for sexual orientation), and CDC 2021 BMI statistics (for weight status). To bridge the gap between the dataset’s fine-grained identity descriptors (e.g., “Buddhist”, “Afro-Latine”) and high-level reference categories (e.g., non-Christian, Black), we implemented semantic mapping dictionaries across multiple axes, including race/ethnicity, religion, gender/sex, and sexual orientation. This approach enabled aggregation of 14,346 religion-related samples, 1,594 gender samples, and 8,550 race/ethnicity samples into population-level categories for KL divergence computation. Full mapping rules are listed in Appendix B.4.1.1.
@@
Annotation Bias. Because HolisticBias itself contains no ground‑truth annotations, any bias arises only after a labeling function $g_{\theta}(x)$ is applied. Following the unified definition in §2.3, we therefore estimate annotation bias by comparing the expected outputs of external scorers across protected groups (Equation 2). We probe the dataset with four commonly used instruments: Gender Polarity, VADER Sentiment, RoBERTa‑Regard, and Perspective‑API Toxicity. We compute group‑level statistics to reveal systematic disparities. Only the toxicity-related graphs are shown here; all other metric plots are included in Appendix B.4.
  • (a)
    Gender Polarity. The Gender Polarity metric (see Appendix B.4) measures token-level association of words with a male/female polarity axis.
    • The heatmap (Figure 18) further reveals specific descriptors aligned with male or female polarity.
@@
  • (b)
    Sentiment. The VADER Sentiment metric evaluates sentence-level polarity. Axis-level differences are minimal (Kruskal–Wallis $p=0.48$), and descriptor-level variation (Appendix B.4) is often tied to emotionally charged words (e.g., “curvy”, “bankrupt”).
  • (c)
    Regard. RoBERTa-based Regard scores (Appendix B.4) highlight positive/negative connotations. Despite uniformly high medians, individual terms like “drug-addicted” (low) and “Christian” (high) introduce descriptor-level variance.
  • (d)
    Toxicity and Sub-Dimensions. We evaluate annotation bias (auxiliary scoring metrics) in the HolisticBias dataset using the Perspective API, which provides six sub-dimensions: TOXICITY, SEVERE_TOXICITY, IDENTITY_ATTACK, INSULT, PROFANITY, and THREAT. These metrics enable us to capture nuanced differences in how identity-related prompts are perceived in terms of harmful or offensive content.
@@
    This pattern persists across sub-dimensions as well. IDENTITY_ATTACK scores often co-occur with TOXICITY patterns, particularly for racial and religious descriptors (see Appendix B.4, Figures 55 and 56). Likewise, SEVERE_TOXICITY and THREAT scores are relatively low overall but spike for marginalized groups when prompts contain specific identity terms or stigmatized descriptors. INSULT scores also align closely with descriptors related to disability, age, or appearance, indicating a tendency for models to associate these identities with ridicule or demeaning language (see Appendix B.4.3.4, Figures 57 and 58).
@@
Across all four instruments, HolisticBias reveals that identity‑conditioned label shifts persist even when the underlying prompt is syntactically neutral. The magnitude of $\hat{B}_{\text{ann}}$ varies by metric and axis but consistently concentrates on historically marginalized descriptors. Consequently, any downstream evaluation that relies on these scorers, by extension, on the labels they generate, must account for this annotation bias to avoid spurious conclusions about model fairness.
Therefore, HolisticBias is not suitable for stereotype leakage analysis though PMI or MI, as it lacks the spontaneous co-occurrence structure necessary for these metrics. This contrasts with naturalistic corpora such as Wikipedia or Reddit, where co-occurrence patterns are organically present and stereotype leakage can be meaningfully quantified. HolisticBias’s strength lies in its ability to test axis-specific model behavior under tightly controlled inputs, not in revealing embedded societal biases within its own textual structure.
3.4 Classification-based Bias Datasets
This subsection covers datasets designed to test bias in classification systems, particularly models that classify text into categories such as sentiment, entity type, or other attributes. These datasets use carefully controlled inputs where demographic features (gender or race) are swapped to evaluate if the model’s predictions are influenced by these attributes. Datasets such as EEC and Bias NLI assess whether a model’s output changes when a sentence’s demographic components, such as the gender of a subject or the race of a person, are modified. These tests often examine if models are more likely to predict a particular label based on demographic information present, testing for bias in tasks such as sentiment analysis, natural language inference (NLI), and entity classification.
@@ 3.4.1 Equity Evaluation Corpus (EEC)
Description. The Equity Evaluation Corpus (EEC), introduced by Kiritchenko and Mohammad (2018), is a benchmark dataset designed to detect demographic bias, specifically race and gender bias, in sentiment analysis systems. It consists of sentence templates with controlled lexical substitutions where only a race- or gender-associated term (e.g., names, pronouns) is varied, enabling a counterfactual comparison of sentiment intensity scores. The goal is to identify whether the systems yield different sentiment predictions for sentences that are semantically identical except for demographic references.
Dataset Taxonomy. EEC is classified as a constrained, form evaluation dataset within the classification, based bias category. It employs a minimal-pair counterfactual design using emotion-laden sentence templates, where variants are created by swapping gender- or race-associated terms, enabling isolated evaluation of bias effects in sentiment analysis systems. The dataset tests whether models yield different sentiment predictions for sentences that are semantically identical except for demographic references, providing a controlled framework for detecting demographic bias in text classification tasks.
In terms of source, EEC employs template-based construction with emotion-laden contexts. The dataset uses a minimal-pair counterfactual design where sentence variants are created by swapping gender- or race-associated words, providing precise control over sentence variability and demographic terms while allowing for targeted fairness evaluations.
The dataset exhibits monolingual linguistic coverage, focusing exclusively on English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, the EEC primarily targets gender and race bias, representing demographic characteristic biases. The dataset focuses on sentiment-related biases in text classification systems, specifically examining how emotional expressions are perceived differently across demographic groups.
Finally, EEC demonstrates high accessibility as a publicly available dataset, likely distributed under a permissive license. This open availability promotes transparency, collaboration, and replicability in fairness research, enabling broad community participation in bias evaluation and mitigation efforts.
Intrinsic Characteristics. The dataset comprises 8,640 English sentences generated from a relatively small number of core templates. Each sentence includes emotional expressions, and the variants differ only in the demographic term used. The text-based format supports standardized input to sentiment analysis models, making it straightforward for quantitative bias evaluations.
@@
However, the dataset has notable limitations. It is highly sensitive to minor rewordings of template structures, raising concerns about robustness. Its reliance on binary gender terms and a small number of templates limits both its inclusivity and linguistic diversity. Moreover, the template-based nature of the data leads to concerns about artificiality and limited ecological validity. Even dataset creators acknowledge that EEC should be viewed as a complementary tool rather than a comprehensive benchmark for bias.
Bias Analysis.
@@
  • Gender. The corpus is split exactly 50 % male / 50 % female. Using 2020 U.S. Census figures (49.1 % male / 50.9 % female) as the reference, the KL divergence is negligible ($D_{\mathrm{KL}}{=}0.0003$), confirming that gender representativeness is virtually perfect.
  • Race. EEC contains only two racial categories and forces them to parity (50 % African-American, 50 % European). Relative to BLS 2024 workforce proportions (13.6 % African-American, 58.9 % European), the divergence is substantial ($D_{\mathrm{KL}}{=}0.2475$). In practical terms, African-American mentions are $\approx\!3.7\times$ over- represented, European mentions are under-represented, and all other racial groups are absent. The corpus therefore serves as a bias stress test rather than a demographic mirror.
Annotation Bias. Because every EEC sentence inherits its “gold” emotion label deterministically from the template itself, the corpus carries no intrinsic annotation bias: for any protected attribute values $a_{1},a_{2}\!\in\!\mathcal{A}$, the empirical estimate in Equation (2) satisfies $\hat{B}_{\mathrm{ann}}\!=\!0$. In practice, however, EEC is fed to external labeling functions $g_{\theta}$—sentiment, toxicity, regard, or polarity scorers, to diagnose whether those instruments exhibit group, conditioned disparities. When we apply widely used scorers to the template pairs, systematic gaps emerge that now qualify as annotation bias under the merged definition:
@@
    Gender Polarity. Figure 21 shows that while unigram scores remain near-zero, indicating no explicit word-level gender bias, both the gender_max_score and gender_wavg_score show noticeable differences. Female-associated samples tend to have significantly higher polarity values than male-associated ones. For example, across emotions, sadness and joy show slightly higher polarity variance compared to anger and fear. Race-wise, European-associated samples show greater polarity spread than African-American counterparts. These results suggest subtle alignment differences between generated text and gendered semantic directions in the embedding space.
@@
    Sentiment. Gender and race also show mild divergence (see Figure 22): female–coded sentences are slightly less negative than male–coded ones ($\Delta\!\approx\!0.04$, d=0.18d\!=\!0.18, $p<.01$); European names score marginally higher than African–American names ($\Delta\!\approx\!0.03$, d=0.14d\!=\!0.14, $p<.05$).
@@
    Regard. Regard scores are fairly consistent across demographic groups (female–male $\Delta\!\approx\!0.02$, d=0.09d\!=\!0.09; Euro–AA $\Delta\!\approx\!0.02$, d=0.08d\!=\!0.08; both $p<.05$) (see Figure 23).
@@
    Toxicity. As expected for a balanced template corpus, overall toxicity scores are low. Emotion is the dominant factor (anger >> fear $\approx$ sadness >> joy), but no reliable demographic gap is detected:
@@
    As shown in Figure 25, similar null results hold for the severe_toxicity, identity_attack, insult, and threat sub-scores (all |d|<0.07|d|\!<\!0.07, p>.10p\!>\!.10), suggesting that Perspective’s classifier does not exhibit annotation bias (auxiliary scoring metrics) under EEC conditions.
@@
    Figure 25: Annotation Bias (Auxiliary Scoring Metrics): Toxicity score distributions (toxicity, severe toxicity, identity attack, insult, threat) by Emotion, Gender, and Race using Perspective API.
Taken together, EEC demonstrates that while the dataset’s native labels are bias‑free, several popular automatic scorers inject annotation bias of varying magnitude—strong for Gender Polarity, moderate for Sentiment, weak for Regard, and negligible for Toxicity. Researchers should therefore normalise or otherwise correct these instruments before attributing unfairness to the models under evaluation.
and the corpus‐level mutual information
$\text{MI}(W_{G};W_{T})\approx 0.$ (34)
@@ 3.4.2 Bias NLI
Description. Bias NLI refers to a collection of datasets and methodological efforts designed to probe bias, particularly gender bias, within the Natural Language Inference (NLI) task. As a classification task where models predict entailment, contradiction, or neutrality between a premise and a hypothesis, NLI serves as a valuable testbed for investigating whether models make predictions based on logical inference or rely on spurious correlations. Many approaches introduce template-based counterfactuals by varying demographic terms in hypotheses while preserving the ground-truth label. Others analyze existing NLI datasets, such as SNLI, to reveal annotation artifacts or demographic term dependencies that may reinforce social biases.
Dataset Taxonomy. Bias NLI is classified as a constrained-form evaluation dataset within the classification-based bias category. It encompasses a collection of datasets and methodological approaches designed to probe bias within the Natural Language Inference (NLI) task, where models predict entailment, contradiction, or neutrality between premises and hypotheses. The approach introduces template-based counterfactuals by varying demographic terms in hypotheses while preserving ground-truth labels, enabling investigation of whether models make predictions based on logical inference or rely on spurious demographic correlations.
In terms of source, Bias NLI employs a hybrid approach combining template-based construction with transformation of existing corpora. The dataset structure varies from manually curated hypothesis–premise pairs to large-scale template-generated corpora, some exceeding five million examples, providing both experimental control and scale while potentially introducing construction artifacts from template generation.
The dataset exhibits monolingual linguistic coverage, focusing primarily on English. This English-centric approach reflects the dominance of English in NLP research but limits the generalizability of fairness conclusions across diverse linguistic and cultural contexts.
Regarding bias typology, Bias NLI primarily targets gender bias, though race and religion have also been considered, representing demographic characteristic biases. In addition to social bias, these datasets shed light on structural issues such as hypothesis-only bias, contradiction-word bias, and attestation bias, where models depend on memorized propositions rather than performing genuine inference, representing dataset construction biases.
Finally, Bias NLI demonstrates variable accessibility, with many datasets publicly released alongside research publications, while others may have restricted access. This mixed accessibility reflects the evolving nature of the field and varying institutional policies regarding dataset release.
@@
Strengths and Limitations. A major strength of Bias NLI is its ability to isolate whether models are truly performing inference or instead exploiting surface-level cues linked to protected attributes. The diagnostic nature of these datasets allows researchers to pinpoint specific sources of bias, such as hypothesis-only bias or word-overlap effects. However, reliance on template-generated sentences can introduce artificiality and make models overly sensitive to phrasing. In some cases, evaluations using Bias NLI may reflect known NLI dataset artifacts more than the presence of social bias, complicating interpretation. Moreover, template-based construction may exacerbate attestation bias, causing models to memorize and reuse stereotypical associations. Finally, while these datasets are effective in highlighting known issues, they may underexplore other forms of latent bias or intersectional effects.
@@
Representativeness Bias. We concatenated the train, dev, and test splits of SNLI (N=570,152N{=}570{,}152 premise–hypothesis pairs) and extracted gender information by keyword matching over both fields. [^9] Pronouns and role nouns such as he, she, man, woman, father, mother, boy, girl, etc. were used. Gender could be detected in 434,367434{,}367 instances ($76.2\%$ coverage). The resulting empirical distribution is
@@
whereas the 2020 U.S. Census baseline is
@@
Applying the KL‐divergence metric defined in Section  2.3,
@@
reveals a substantial $27$-percentage-point over-representation of masculine references. This skew implies that fairness evaluations based on BiasNLI may inflate performance on male-coded language while offering weaker evidence of model behavior on female-coded inputs, underscoring the need for rebalancing or stratified reporting.
Annotation Bias. Under the merged definition in §2.3, annotation bias subsumes both human‑base and model‑based labeling disparities. For Bias NLI the labeling function $g_{\theta}$ is deterministic: every protected attribute value appears in the same set of template slots and receives the same three‑way label distribution by design. Consequently, the theoretical expectation in Equation (2) is
$\mathbb{E}[g_{\theta}(x)\mid A=a_{1}]\;=\;\mathbb{E}[g_{\theta}(x)\mid A=a_{2}]\quad\forall\,a_{1},a_{2}\in\mathcal{A},$
yielding $B_{\text{ann}}=0$ absent implementation errors. In practice, minor deviations can arise if (i) certain demographic terms appear in fewer templates, or (ii) the corpus is later scored with an external metric (e.g., Perspective API) whose outputs vary across groups. However, such metric‑driven disparities would speak to model bias in the auxiliary scorer, not to the inherent labels of Bias NLI itself. Because the dataset lacks gold demographic annotations and any group membership must be inferred heuristically from surface terms, aggregate metric comparisons would be noisy and risk conflating lexical mention with true identity. We therefore conclude that Bias NLI is intentionally annotation‑bias‑neutral and ill‑suited for differential group analyses without additional, verified demographic metadata.
Stereotype Leakage. To measure how demographic references co-occur with socially stereotyped traits, we apply the PMI and MI estimators introduced in Section 2. We use fixed group and trait vocabularies, covering gender, race, religion, age, and disability markers, as well as occupation-, personality-, behavior-, and social-status–related traits, specified in Appendix A.2 and used throughout the analysis. For each sentence in the BiasNLI corpus, we slide a token window of size 5 and count co-occurrences of $(w_{g},w_{t})$ pairs. We then smooth and normalize the resulting co-occurrence probabilities to compute pairwise PMI and corpus-level MI.
@@
Figure 26 illustrates the top 10 most stereotypical co-occurrence pairs according to PMI. Among the highest are (jewish, rich), (black, unsuccessful), (him, aggressive), and (african, poverty). These combinations suggest stereotypical associations between religion and wealth, race and socioeconomic failure, and gender and emotional dominance—aligning with historically attested social biases.